# Degrees of Freedom for Radiating Systems

### v1

@@ IV NDoF and shadow area
Determination of NDoF from the number of radiation modes (11) greater than unity or from the effective NDoF (13) is easily achieved numerically for arbitrary shaped objects and material parameters [8], see Fig. 7. This numerical evaluation is complemented by an analytical examination, providing valuable physical insight and understanding. Specifically, it demonstrates that the radiating NDoF in the electrically large limit is proportional to the shadow area measured in squared wavelengths [10].
@@
$\max A_{\mathrm{eff}}=\lambda^{2}\mathbf{F}^{\text{H}}\mathbf{R}^{-1}\mathbf{F}=\frac{\lambda^{2}}{16\pi^{2}}\mathbf{a}^{\text{H}}\mathbf{U}\mathbf{R}^{-1}\mathbf{U}^{\text{T}}\mathbf{a},$ (15)
where the field is expand in spherical waves with expansion coefficients collected in a column matrix $\mathbf{a}$ using $\mathbf{F}=\mathbf{U}^{\text{T}}\mathbf{a}/(4\pi)$ [27]. Diagonalizing the optimization problem (14) using radiation modes (11) induces a change of variables $\tilde{\mathbf{a}}=\mathbf{Q}\mathbf{a}$ with a unitary matrix $\mathbf{Q}$ having elements $Q_{nm}$
with $\sum_{m}Q^{2}_{nm}=1$. Taking the average of $\max A_{\mathrm{eff}}$ over all directions and polarizations using the plane wave expansion $a_{n}=4\pi\mathrm{j}^{\tau-1-l}\hat{\boldsymbol{e}}\cdot\boldsymbol{Y}_{n}(\hat{\boldsymbol{k}})$ [18] with spherical harmonics $\boldsymbol{Y}_{n}$ and $\hat{\boldsymbol{e}}\cdot\hat{\boldsymbol{k}}=0$ results in
and using that $\hat{\boldsymbol{k}}\cdot\boldsymbol{Y}(\hat{\boldsymbol{k}})=0$ extends (18) to an identity dyadic. The same relation (17) holds for the orthogonal combination $\langle|\tilde{a}_{n}|^{2}\rangle=\sum Q^{2}_{mn}\langle|a_{n}|^{2}\rangle=2\pi$ giving a simple identity for the average maximal effective area expressed in radiation modes
@@
$\nu_{n}=\frac{\varrho_{n}}{1+\varrho_{m}}\approx\begin{cases}1&n<N_{\mathrm{A}}\\ 0&n>N_{\mathrm{A}},\end{cases}$ (21)
where $N_{\mathrm{A}}$ denotes the NDoF for the given shape and frequency. The distribution is continuous, which means that a few modes have efficiencies between 1 and 0. The transition from large to small modal efficiencies decreases with increasing electrical size, see Fig. 5.
@@
$\langle\max A_{\mathrm{eff}}\rangle\approx\frac{\lambda^{2}}{8\pi}N_{\mathrm{A}}$ (22)
Rewriting the affective area in the partial gain, $G=4\pi A_{\mathrm{eff}}/\lambda^{2}$, suggest that the NDoF can alternatively be expressed as
$N_{\mathrm{A}}\approx 2\langle\max G\rangle=\sum\frac{\varrho_{n}}{1+\varrho_{n}}$ (23)
@@
The radiation modes are related to the geometrical structure by using that the maximal effective area in a direction $\hat{\boldsymbol{r}}$ approach the geometrical cross section, $A_{\mathrm{s}}(\hat{\boldsymbol{r}})$, in the electrically large limit [31] $\max A_{\mathrm{eff}}(\hat{\boldsymbol{r}})\to A_{\mathrm{s}}(\hat{\boldsymbol{r}})$ and similarly for the average
$\langle\max A_{\mathrm{eff}}\rangle\to\langle A_{\mathrm{s}}\rangle\quad\text{as }ka\to\infty.$ (24)
This connects the radiation modes and geometrical properties of the region $\varOmega$ with the maximal effective area (24) expressed solely in geometrical parameters producing the asymptotic NDoF
$N_{\mathrm{A}}\approx\frac{8\pi\langle A_{\mathrm{s}}\rangle}{\lambda^{2}}=\frac{2|A_{\mathrm{s}}|}{\lambda^{2}}\quad\text{as }ka\to\infty,$ (25)
@@
Radiation modes for six different objects are depicted in Fig. 8. The radiation mode eiganvalues $\varrho_{n}$ are normalized with the surface resistivity $R_{\mathrm{s}}$, and the mode index $n$ is scaled with the NDoF, $N_{\mathrm{A}}$, based on the shadow area (25), as shown in Tab. I. The $N_{\mathrm{A}}$ dominant normalized radiation modes have approximately unit magnitude, after which they rapidly decay.
Figure 8: Normalized radiation modes $\varrho_{n}$ for six objects of the electrical size $ka=50$ plotted versus the normalized mode index $n/N_{\mathrm{A}}$ (25).
The average shadow area $\langle A_{\mathrm{s}}\rangle$ can be determined analytically or numerically. For convex shapes it reduces to a quarter of the surface area $A$ [32]
@@
This can be derived from the fact that each ray intersecting the object intersects the boundary twice [32], as shown in Fig. 9. For non-convex objects, rays might intersect the boundary more than twice. This can be interpreted as shadowing of some regions, or equivalently, that some boundary points cannot radiate undestructively over $2\pi$ steradian to the farfield. For convex objects, the NDoF (25) simplifies as
$N_{\mathrm{A}}\approx\frac{8\pi\langle A_{\mathrm{s}}\rangle}{\lambda^{2}}\stackrel{{\scriptstyle\begin{subarray}{c}\mathrm{convex}\\ \mathrm{object}\end{subarray}}}{{=}}\frac{2\pi A}{\lambda^{2}}=\frac{k^{2}A}{2\pi}\quad\text{as }ka\in\infty.$ (27)
@@
Figure 9: DoF for a region $\varOmega\in\mathbb{R}{}^{3}$. (left) convex object having rays intersecting the boundary $\partial\varOmega$ twice. (right) non-convex object with some rays intersecting the boundary more than twice.
@@
Shadow areas for the objects in Tab. I are independent of the azimuthal coordinate $\phi$ and plotted as function of the polar angle $\theta$ in Fig. 10. The sphere has $A_{\mathrm{s}}(\theta)=\pi a^{2}$ for all directions. The disc is also approximately $A_{\mathrm{s}}(0)\approx\pi a^{2}$ in the normal direction but much smaller seen from the side $A_{\mathrm{s}}(\theta)\approx\pi a^{2}\cos\theta$. The solid cylinder and connected discs coincide for illuminations from above, $\theta\in[0,0.2\pi]$, but start to differ for illuminations from the side. The opposite holds for the cylinder compared with the open cylinder where the opening is observed from the above but not from the side. The corrugated cylinder has similar cross sectional area as the cylinder but differ for illuminations from the side where the corrugations are visible. The difference at $\theta=0$ is due to the different aspect ratio, see Tab. I. The shadow areas are bounded by $\pi a^{2}$ which is reached for a sphere for all directions but also approximately for thin discs. The minimal values can approach zero as observed for the open cylinder at $\theta=0$ and disc at $\theta=\pi/2$. The NDoF depends on the total shadow area (25), which is determined by integration of the curves in Fig. 10 weighted by $\sin\theta$.
The NDoF (25) depends on the material properties such as surface resistivity $R_{\mathrm{s}}$ or bulk resistivity $\rho_{\mathrm{r}}$. Metals typically have values $R_{\mathrm{s}}/\eta_{0}$ and $k\rho_{\mathrm{r}}/\eta_{0}$ below $10^{-4}$ producing a distinct difference between efficient and inefficient modes as exemplified in Fig. 8 and Fig. 5. Highly resistive materials have much larger resistivity giving fewer efficient modes and hence less DoF. Materials have here for simplicity been treated as homogeneous but the formulation applies equally well for inhomogeneous regions.
@@
$\int_{\varOmega}|\boldsymbol{J}(\boldsymbol{r})|^{2}\operatorname{dV}\!\approx\mathbf{I}^{\text{H}}\mathbf{\Psi}\mathbf{I}$ (28)
@@ V Electric and magnetic currents
Allowing for magnetic currents increases the NDoF compared with the solely electric case. This is particularly evident for electrical sizes where the geometrical structure is not resolved by the wavelength. NDoF of infinitely thin sheets such as planar regions do not directly follow from Weyl’s law as they do not have an inner region. Considering e.g., a planar region with only electric currents enforces a symmetry of the radiated field in the up and down directions reducing the NDoF [8]. This symmetry is broken by allowing both electric and magnetic currents on the thin sheet and hence making the structure behave as a volumetric region.
@@
Considering first the case with both electric, $\boldsymbol{J}$, and magnetic currents, $\boldsymbol{M}$, shown by solid curves and markers in Fig. 11. The radiation modes for oblate spheroids with $\xi=0.01$ and $\xi=0$ (infinitely thin disc) overlap. This is due to the negligible thickness of the $\xi=0.01$ oblate spheroid. The radiation modes $\varrho_{n}$ have a cutoff around the NDoF estimate (25) based on the average shadow area $\langle A_{\mathrm{s}}\rangle=\pi a^{2}/2+\pi a^{2}\xi^{2}/(4e)\ln((1+e)/(1-e))$ with $e^{2}=1-\xi^{2}$, which approaches $\pi a^{2}/2$ as $\xi\to 0$. The slope of $\varrho_{n}$ for $n>N_{\mathrm{A}}$ becomes steeper as $ka$ increases, similar to the results for the sphere in Fig. 5. Here we also note that the disc $\xi=0$ is treated as having a top and bottom, which can equivalently be treated as half the resistivity for the disc.
@@
Needle shaped prolate spheroidals are similarly one dimensional showing approximately a different scaling as evident from Weyl’s law (4) before their thickness is resolved.
@@
The presented derivation of the NDoFs remain valid for non-connected and multiple objects, as demonstrated by the two square plates in Figs. 12 and 13. Fig. 12 illustrates the radiation modes $\varrho_{n}$ for two plates separated by a distance $d/\ell\in\{0.01,0.1,1,10\}$. The modal index $n$ is normalized with the asymptotic NDoF $N_{\mathrm{A}}$ based on the average shadow area $\langle A_{\mathrm{s}}\rangle$. Notably, the radiation modes $\varrho_{n}$ exhibit a cutoff around $N_{\mathrm{A}}$. The shadow area can be numerically evaluated by considering the projected area of the two squares in a plane parallel to the plates. This resulting shadow area, plotted versus the distance $d$, is illustrated in Fig. 13 by the curve $\langle A_{\mathrm{s}}(d/\ell)\rangle/\langle A_{\mathrm{s}}(0)\rangle$, where $\langle A_{\mathrm{s}}(0)\rangle=\ell^{2}/2$ represents the average shadow area for a single square, as shown in equation (26). As the distance increases, the average shadow area also increases and approaches $2\langle A_{\mathrm{s}}(0)\rangle$ as $d/\ell\to\infty$, corresponding to the NDoF of a plate with an area of $2\ell^{2}$.
@@
VI Inverse source problems
$\varOmega$${\boldsymbol{J}},\boldsymbol{M}$$\mathbf{f}$Rxb)
Figure 14: Inverse source problem for a region with surface $\varOmega$ and equivalent (surface) currents $\boldsymbol{J}({\boldsymbol{r}}),\boldsymbol{M}({\boldsymbol{r}})$ with measurements in the far field or on a circumscribing sphere.
The concept of NDoF for radiating systems, as analyzed in Sec. IV, also applies to inverse source problems. Consider a region composed of some linear material. By employing the equivalence principle, the radiated field can be represented by electric $\boldsymbol{J}$ and magnetic $\boldsymbol{M}$ surface current densities on a surface $\varOmega$ surrounding the region [33]. For simplicity, we let this surface coincide with with boundary $\varOmega$ of the region. With all sources inside $\varOmega$, Love’s theorem or the null field condition can be used to relate the electric and magnetic currents in the inverse source problem [34, 35].

### v2

@@ IV NDoF and shadow area
Determination of NDoF from the number of radiation modes (11) greater than unity $N_{1}$ or from the effective NDoF (13) $N_{\mathrm{e}}$ is easily achieved numerically for arbitrary shaped objects and material parameters [8], see Fig. 7. This numerical evaluation is complemented by an analytical examination, providing valuable physical insight and understanding. Specifically, it demonstrates that the radiating NDoF in the electrically large limit is proportional to the shadow area measured in squared wavelengths [10].
@@
$\max A_{\mathrm{eff}}(\hat{\boldsymbol{k}},\hat{\boldsymbol{e}})=\lambda^{2}\mathbf{F}^{\text{H}}\mathbf{R}^{-1}\mathbf{F}$ (15)
and taking the average of $\max A_{\mathrm{eff}}(\hat{\boldsymbol{k}},\hat{\boldsymbol{e}})$ over all directions $\hat{\boldsymbol{k}}$ and polarizations $\hat{\boldsymbol{e}}$
$\langle\max A_{\mathrm{eff}}\rangle=\frac{1}{8\pi^{2}}\int_{2\pi}\int_{4\pi}\max A_{\mathrm{eff}}(\hat{\boldsymbol{k}},\hat{\boldsymbol{e}})\operatorname{d}\!\Omega_{\hat{\boldsymbol{k}}}\operatorname{d}\!\Omega_{\hat{\boldsymbol{e}}}$ (16)
and similarly for the average maximal partial gain $\langle\max G\rangle=4\pi\langle\max A_{\mathrm{eff}}\rangle/\lambda^{2}$. By diagonalizing the optimization problem (14) using the radiation modes (11) and performing the directional and polarization averaging (16) detailed in App. B, we derive a straightforward expression for the average maximal effective area, formulated in terms of the radiation modes
@@
$\nu_{n}=\frac{\varrho_{n}}{1+\varrho_{n}}\approx\begin{cases}1&n<N_{1}\\ 0&n>N_{1},\end{cases}$ (19)
where $N_{1}$ denotes the NDoF for the given shape and frequency. The distribution is continuous, which means that a few modes have efficiencies between 1 and 0. The transition from large to small modal efficiencies decreases with increasing electrical size, see Fig. 5.
@@
$\langle\max A_{\mathrm{eff}}\rangle\approx\frac{\lambda^{2}}{8\pi}N_{1}.$ (20)
Rewriting the partial effective area in the partial gain, $G=4\pi A_{\mathrm{eff}}/\lambda^{2}$, suggest that the NDoF can alternatively be expressed as
$N_{1}\approx 2\langle\max G\rangle=\frac{1}{2}\sum\frac{\varrho_{n}}{1+\varrho_{n}},$ (21)
@@
The radiation modes are related to the geometrical structure by using that the maximal effective area in a direction $\hat{\boldsymbol{k}}$ approaches the shadow area (geometrical cross section for convex shapes), $A_{\mathrm{s}}(\hat{\boldsymbol{k}})$, in the electrically large limit [36] $\max A_{\mathrm{eff}}(\hat{\boldsymbol{k}})\to A_{\mathrm{s}}(\hat{\boldsymbol{k}})$ and similarly for the average over polarizations and directions
$\langle\max A_{\mathrm{eff}}\rangle\to\langle A_{\mathrm{s}}\rangle=\frac{1}{4\pi}\int_{4\pi}A_{\mathrm{s}}(\hat{\boldsymbol{k}})\operatorname{d}\!\Omega_{\hat{\boldsymbol{k}}}$ (22)
as $ka\to\infty$. This connects the radiation modes and geometrical properties of the region $\varOmega$ with the maximal effective area (22) expressed solely in geometrical parameters producing the asymptotic NDoF
$N_{1}\approx N_{\mathrm{A}}=\frac{8\pi\langle A_{\mathrm{s}}\rangle}{\lambda^{2}}=\frac{2|A_{\mathrm{s}}|}{\lambda^{2}}\quad\text{as }ka\to\infty,$ (23)
@@
[…]
Figure 8: Illustration of the shadow area $A_{\mathrm{s}}(\hat{\boldsymbol{k}})$ of a tilted corrugated cylinder for illumination directions $\hat{\boldsymbol{x}}$, $\hat{\boldsymbol{y}}$, and $\hat{\boldsymbol{z}}$, as projected onto a plane perpendicular to each respective illumination direction.
The shadow area, $A_{\mathrm{s}}(\hat{\boldsymbol{k}})$, is determined by illuminating the (opaque) region in the direction $\hat{\boldsymbol{k}}$ and projecting the resulting shadow onto a plane perpendicular to this direction, as shown in Fig. 8. For convex shapes, the shadow area is equivalent to the maximum cross-sectional area in the given direction. Note that the shadow area is symmetric $A_{\mathrm{s}}(\hat{\boldsymbol{k}})=A_{\mathrm{s}}(-\hat{\boldsymbol{k}})$.
The NDoF and shadow area are first analyzed for a spherical region with radius $a$. Due to the spherical symmetry, the shadow area is independent of the illumination direction, $A_{\mathrm{s}}(\hat{\boldsymbol{k}})=\pi a^{2}$, corresponding to the cross-sectional area. Consequently, the average shadow area (22) over all directions is also $\langle A_{\mathrm{s}}\rangle=\pi a^{2}$. The NDoF, given by (23), is $2(ka)^{2}$, consistent with the classical result in (6). The corresponding average maximal gain, derived from (21), is $(ka)^{2}$, which matches the leading term in Harrington’s gain estimate [37].
For a spherical region, the total surface area, $A=4\pi a^{2}$, is exactly four times the average shadow area, $A=4\langle A_{\mathrm{s}}\rangle$. This relationship is a general result for convex shapes, originally derived by Cauchy [38]
@@
This identity follows from the principle that each ray intersecting the object also intersects its boundary twice [38], as illustrated in Fig. 9. For non-convex objects, rays might intersect the boundary more than twice. This can be interpreted as shadowing of some regions, or equivalently, that some boundary points cannot radiate undestructively over $2\pi$ steradian to the farfield. For convex objects with surface area $A$, the NDoF (23) reduces to
$N_{\mathrm{A}}\approx\frac{8\pi\langle A_{\mathrm{s}}\rangle}{\lambda^{2}}\stackrel{{\scriptstyle\begin{subarray}{c}\mathrm{convex}\\ \mathrm{object}\end{subarray}}}{{=}}\frac{2\pi A}{\lambda^{2}}=\frac{k^{2}A}{2\pi}\quad\text{as }ka\to\infty.$ (25)
@@
Figure 9: Shadow area for a region $\varOmega\in\mathbb{R}{}^{3}$. (left) convex object having rays intersecting the boundary $\partial\varOmega$ twice. (right) non-convex object with some rays intersecting the boundary more than twice.
Radiation modes for six different objects are depicted in Fig. 10. The radiation mode eigenvalues $\varrho_{n}$ are normalized with the surface resistivity $R_{\mathrm{s}}$, and the mode index $n$ is scaled with the NDoF, $N_{\mathrm{A}}$, based on the shadow area (23), as shown in Tab. I. The $N_{\mathrm{A}}$ dominant normalized radiation modes have approximately unit magnitude, after which they rapidly decay.
Figure 10: Normalized radiation modes $\varrho_{n}$ for the six objects in Tab. I of the electrical size $ka=50$ plotted versus the normalized mode index $n/N_{\mathrm{A}}$ (23).
@@
Shadow areas for the objects in Tab. I are independent of the azimuthal coordinate $\phi$ and plotted as function of the polar angle $\theta$ in Fig. 11. The sphere has $A_{\mathrm{s}}(\theta)=\pi a^{2}$ for all directions. The disc is also approximately $A_{\mathrm{s}}(0)\approx\pi a^{2}$ in the normal direction but much smaller seen from the side $A_{\mathrm{s}}(\theta)\approx\pi a^{2}\cos\theta$. The solid cylinder and connected discs coincide for illuminations from above, $\theta\in[0,0.2\pi]$, but start to differ for illuminations from the side. The opposite holds for the cylinder compared with the open cylinder, where the opening is observed from the above but not from the side. The corrugated cylinder has similar cross-sectional area as the cylinder but differ for illuminations from the side where the corrugations are visible, cf.,  Fig. 8. The difference at $\theta=0$ is due to the different aspect ratio, see Tab. I. The shadow areas are bounded by $\pi a^{2}$ which is reached for a sphere for all directions but also approximately for thin discs. The minimal values can approach zero, as observed for the open cylinder at $\theta=0$ and disc at $\theta=\pi/2$. The NDoF depends on the total shadow area (23), which is determined by integration of the curves in Fig. 11 weighted by $\sin\theta$.
The NDoF (23) depends on the material properties such as surface resistivity $R_{\mathrm{s}}$ or bulk resistivity $\rho_{\mathrm{r}}$. Metals typically have values $R_{\mathrm{s}}/\eta_{0}$ and $k\rho_{\mathrm{r}}/\eta_{0}$ below $10^{-4}$ producing a distinct difference between efficient and inefficient modes as exemplified in Fig. 10 and Fig. 5. Highly resistive materials have much larger resistivity, giving fewer efficient modes and hence less DoF. Materials have here for simplicity been treated as homogeneous, but the formulation applies equally well for inhomogeneous regions.
@@
$\int_{\varOmega}|\boldsymbol{J}(\boldsymbol{r})|^{2}\operatorname{dV}\!\approx\mathbf{I}^{\text{H}}\mathbf{\Psi}\mathbf{I}.$ (26)
@@ V Electric and magnetic currents
Allowing for magnetic currents increases the NDoF compared with the solely electric case. This is particularly evident for electrical sizes where the geometrical structure is not resolved by the wavelength. NDoF of infinitely thin sheets such as planar regions do not directly follow from Weyl’s law, as they do not have an inner region. Considering e.g., a planar region with only electric currents enforces a symmetry of the radiated field in the up and down directions reducing the NDoF [8]. This symmetry is broken by allowing both electric and magnetic currents on the thin sheet, and hence making the structure behave as a volumetric region.
@@
Considering first the case with both electric, $\boldsymbol{J}$, and magnetic currents, $\boldsymbol{M}$, shown by solid curves and markers in Fig. 12. The radiation modes for oblate spheroids with $\xi=0.01$ and $\xi=0$ (infinitely thin disc) overlap. This is due to the negligible thickness of the $\xi=0.01$ oblate spheroid. The radiation modes $\varrho_{n}$ have a cutoff around the NDoF estimate (23) based on the average shadow area $\langle A_{\mathrm{s}}\rangle=\pi a^{2}/2+\pi a^{2}\xi^{2}/(4e)\ln((1+e)/(1-e))$ with $e^{2}=1-\xi^{2}$, which approaches $\pi a^{2}/2$ as $\xi\to 0$. The slope of $\varrho_{n}$ for $n>N_{\mathrm{A}}$ becomes steeper as $ka$ increases, similar to the results for the sphere in Fig. 5. Here, we also note that the disc $\xi=0$ is treated as having a top and bottom, which is equivalent to using half the resistivity for the disc.
@@
Needle shaped prolate spheroidals are similarly one dimensional, showing approximately a different scaling as evident from Weyl’s law (4) before their thickness is resolved.
@@
The presented derivation of the NDoFs remain valid for non-connected and multiple objects, as demonstrated by the two square plates in Figs. 13 and 14. Fig. 13 illustrates the radiation modes $\varrho_{n}$ for two plates separated by a distance $d/\ell\in\{0.01,0.1,1,10\}$. The modal index $n$ is normalized with the asymptotic NDoF $N_{\mathrm{A}}$ based on the average shadow area $\langle A_{\mathrm{s}}\rangle$. Notably, the radiation modes $\varrho_{n}$ exhibit a cutoff around $N_{\mathrm{A}}$. The shadow area can be numerically evaluated by considering the projected area of the two squares in a plane parallel to the plates. This resulting shadow area, plotted versus the distance $d$, is illustrated in Fig. 14 by the curve $\langle A_{\mathrm{s}}(d/\ell)\rangle/\langle A_{\mathrm{s}}(0)\rangle$, where $\langle A_{\mathrm{s}}(0)\rangle=\ell^{2}/2$ represents the average shadow area for a single square, as shown in equation (24). As the distance increases, the average shadow area also increases and approaches $2\langle A_{\mathrm{s}}(0)\rangle$ as $d/\ell\to\infty$, corresponding to the NDoF of a plate with an area of $2\ell^{2}$.
@@
VI Inverse source problem
$\varOmega$${\boldsymbol{J}}({\boldsymbol{r}})$$\mathbf{f}$Rxb)
Figure 15: Inverse source problem for a region with surface $\varOmega$ with measurements in the far field or on a circumscribing sphere.
The concept of NDoF for radiating systems, as analyzed in Sec. IV, also applies to inverse source problems. Consider a region composed of some linear material. By employing the equivalence principle, the radiated field can be represented by electric $\boldsymbol{J}$ and magnetic $\boldsymbol{M}$ surface current densities on a surface $\varOmega$ surrounding the region [39]. For simplicity, we let this surface coincide with the boundary $\varOmega$ of the region. With all sources inside $\varOmega$, Love’s theorem or the null field condition can be used to relate the electric and magnetic currents in the inverse source problem [40, 41].